OpenStudy (anonymous):

Jennifer is twice as old as Ramon. The sum of their ages 7 yrs ago was 13. How old are they now?

OpenStudy (saifoo.khan):

R=X J=2X (X-7)+(2X-7)=13 X-7+2X-7=13 3X=13+14 3X=27 X=27/3 X=9 YEARS OLD IS RAMON NOW. 2*9=18 YEARS OLD IS JENNIFER NOW.